# Higher-Dimensional MOG dark compact object: shadow behaviour in the light of EHT observations

### v1

@@ III Shadow behaviour of the higher-dimensional MOG dark compact object
A part of the light that would reach the observer is deflected by the black hole’s gravitational field when it passes in front of it. The apparent shape of the black hole is the boundary of the shadow, which is created when photons fall into the black hole and form a dark area known as the shadow. The general formulae needed to determine the deflection angle, energy emission rate, and shadow shape for the higher-dimensional general ansatz (8)—which requires examining a test particle’s motion in spacetime—are provided in this section. Then we apply higher-dimensional MOG dark compact object line element on the formulas deduced in this section.
In MOG theory with extra dimensions, we analyze the shadow and deflection angle of the dark compact object by taking arbitrary values for the STVG parameter, $\alpha=0,0.2,0.45,0.5$, and $0.674$. Furthermore, we consider that the additional dimensions count $D=4,5,...,11$ (keep in mind that $D=4$ represent one temporal in addition to three spatial dimensions as usaul). In Sau:2022afl , the authors consider regular MOG dark compact objects, using nonlinear field equations to determine the effects of $\phi_{\mu}$ for $\alpha>0.674$ by adopting the ADM mass, here we examine the case for $\alpha\leq\alpha_{crit}$.
III.1 Effective potential
Initially, we aim to examine the effective potential behaviour of the dark compact object with arbitrary dimensions. To do this, we start with the Lagrangian of the test particle that has the following form
@@ III.1 Effective potential
where $i=1,2,3,...,D-3$, and $L$ is the angular momentum and $E$ is the energy of the test particle. We apply the Carter approach to study the geodesic equations and the Hamilton-Jacobi method to analyze photon orbits around the black hole. We need to generalize these methods to higher dimensions for our work. The Hamilton-Jacobi method in extra dimensions has the following form
@@
in which $S$ is the Jacobi action of the test particle. Inserting the general ansatz. (8) with arbitrary dimensions into Eq. (20), we have
@@
where $m$ is the rest mass of the test particle. Since the photon is the test particle used to analyze the shadow behavior of black holes, we set $m=0$. Also by inserting the action (22) into Eq. (21) we derive the following equations
@@
where “$+$” and “--” signs denote the outgoing and ingoing radial directions of the motion of photon, respectively. Also we have the motion of photon in the spacetime as follows which can be derived by Eqs. (26)-(29).
@@
Rewriting the radial null geodesic equation (27) yields the effective potential,which has the following form:
@@
The unstable circular orbits of photons define the boundary of apparent shape of the black hole. They correspond with the maximum value of the effective potential, which occurs at a distance, referred as photon sphere radius $r_{0}$ satisfyig the equations below
@@
Therefore, the least value of the roots of the following equation is the photon sphere radius $r_{0}$ associated with the maximum of the black hole’s effective potential in spacetime (8) with arbitrary dimensions,
@@
For various values of $D$ and $\alpha$, the behavior of the effective potential for the higher dimensional MOG dark compact object is shown in Fig. 2 as a function of radial coordinate $r$. The photon sphere radius $r_{0}$, which corresponds to each value of $D$ and $\alpha$, is where the effective potential in this figure peaks, and as $r\rightarrow\infty$, the effective potential approaches a steady value. As we can observe from Fig. 2a, the higher-dimensional MOG dark compact object’s effective potential grows as $D$ increases for a constant value of $\alpha$. Furthermore, we discover from Figs. 2b and 2c that raising $\alpha$ for $D=4,6$ causes the dark compact object’s effective potential to be amplified. By examining these figures it is obvious that the impact of higher dimensions dominates the effect of $\alpha$ in the MOG dark compact object by increasing the spacetime dimensionality. Figure 2 illustrates how $D$ and $\alpha$ impact the shadow boundary of the MOG dark compact object by increasing the spacetime dimensionality. This is because the location of the black hole’s maximum effective potential, or the photon sphere radius $r_{0}$, describes the black hole’s shadow boundary.
III.2 Geometrical shapes of shadow
@@
Therefore, one can rewrite the effective potential and also, the function $\mathcal{R}$ in terms of these impact parameters as
@@
Finally, by inserting Eq. (36) into Eq. (32) one can find the following equation for two unknowns $\xi$ and $\eta$
@@
Consequently, using Eq. (37), the photon sphere radius obtained from Eq. (33) yields the quantity $\eta+\xi^{2}$. It is evident that the quantity $\eta+\xi^{2}$ has the dimension of the length square, and $r_{0}$ has the dimension of the length.
Usually $\lambda$ and $\psi$ are used as celestial coordinates to describe the geometric shape of the shadow as perceived by the observer’s frame Vazquez:2003zm . These coordinates can be read as follows
@@
We will now demonstrate the geometrical shape of the MOG dark compact object’s shadow in extra dimensions on the observer’s sky using the celestial coordinates that were presented above. To begin with, we should collect some numerical data for the black hole’s related $r_{eh},r_{0}$, and $\sqrt{\eta+\xi^{2}}$. For the higher-dimensional MOG dark compact object, the photon sphere radius may be obtained by inserting Eq. (9) into Eq. (33). Furthermore, the radius of shadow circles for the higher-dimensional MOG dark copact object in celestial coordinates may be found by utilizing Eq. (39) and applying Eq. (9) into Eq. (37). In Table 1 we collect the numerical data associated with $r_{e}h$, and $r_{0}$ for $D=4,5,6,7,8$ and some different values of $\alpha$.
@@
Plotting the shadow circles of the MOG dark compact object with extra dimensions for various values of $\alpha$ and $D$ is possible using the information given in Table 1. Figure 3 shows the geometrical shapes of the MOG dark compact object’s shadow with extra dimensions in celestial coordinates for various values of $\alpha$ and $D$. As one can see it shows plots for fixed values of $\alpha$ and illustrates how the black hole’s shadow shapes get smaller as dimension increases for a certain value of $\alpha$. Therefore, by reducing the size of shadow’s geometrical shape, the extra dimensions have a major impact on the black hole’s shadow.
@@ III.3 Energy Emission rate
Thus, the complete form of the energy emission rate of higher-dimensional black holes can be read as
@@
In order to obtain the values of the Hawking temperature of the MOG dark compact object with extra dimensions for various values of $\alpha$ and $D$, we now use the values of $r_{eh}$ from Table 1. Next, by inserting the values of $\sqrt{\eta+\xi^{2}}$ into Eq. (40), one can determine the values of $\sigma_{lim}$ for various values of $\alpha$ and $D$ which correspond to the higher-dimensional MOG dark compact object. In order to obtain the energy emission rate in terms of various values of $\alpha$ and $D$ associated with the higher-dimensional MOG dark compact object, one can finally use the values of the Hawking temperature and $\sigma_{lim}$ in Eq. (41).
@@
Fig. 5 shows the energy emission rate as a function of the emission frequency for the higher-dimensional MOG dark compact object. The energy emission rate of MOG dark compact object with extra dimensions greatly increases by growing $D$, as shown in Figs. 5a and 5b, at a fixed value of $\alpha$. Consequently, we discover that in STVG with higher dimensions, extra dimensions speed up the MOG compact object’s evaporation. Also, the impact of the STVG parameter is to amplify the energy emission rate of MOG dark compact object, which causes it’s evaporation to accelerate.
@@ III.4 Deflection Angle
We aim to give a framework for examining the deflection angle of higher-dimensional black holes in the line element (8). We intend to apply the Gauss-Bonnet theorem Gibbons:2008rj ; Arakida:2017hrm . To begin, we compute the optical metric on the equatorial hyperplane $\theta_{i}=\pi/2$ using the line element (8). On this hyperplane, we set $d\theta_{D-2}^{2}=d\phi^{2}$ to find
@@
Finally, using the straight light ray approximation $r(\phi)=\xi/\sin\phi$, the Gauss-Bonnet theorem results in the following expression to calculate the deflection angle (for more details, see Refs. Gibbons:2008rj ; Arakida:2017hrm and references therein)
@@
Furthermore, the surface element of the optical metric (42) for the higher-dimensional MOG dark compact object in correspondance with the metric coefficients (9) can be found approximately as follows
@@
Now, employing Eqs. (49) and (50) in the deflection angle formula (48) leads to the deflection angle of the higher-dimensional MOG dark compact object as follows
@@
The behavior of the deflection angle of the higher-dimensional MOG dark compact object is illustrated in Fig. 6 for different values of $D$ with respect to $\alpha=0.45$ in Fig. 6a and for different values of $\alpha$ with respect to $D=4,5$ in Figs. 6b and 6c, respectively. Fig. 7a, shows us that decreasing the value of the impact parameter $\xi$ results in extremely increasing the deflection angle of the MOG dark compact object. From Fig. 6a, we see that for a fixed value of STVG parameter, the deflection angle of the MOG dark compact object reduces by growing the number of extra dimensions. In Fig. 6b and Fig. 6c we see that for a fixed value of $D$, the deflection angle of the black hole decreases by increasing STVG parameter $\alpha$.
IV Constraints from EHT observations of M87*
@@
Figure 7 indicates the behavior of the shadow radius of the higher-dimensional MOG dark compact object in comparison with the EHT’s shadow size of M87* within $1$-$\sigma$ uncertainties given in Eq. (52) versus the STVG parameter $\alpha$. In this figure, the white (unshaded) region denotes the $1$-$\sigma$ confidence level while the brown (shaded) areas are the excluded regions, which are incompatible with the EHT observations associated with shadow radius of M87*. From this figure, we see that the shadows of the higher-dimensional MOG dark compact object associated with $D=5,6,7$ are incompatible with the observations of EHT. However, the shadow of the MOG dark compact object with $D=4$ lies within the $1$-$\sigma$ confidence level of EHT data, so that in the range $0.2790\leq\alpha<$ 0.279 the shadow radius of the four dimensional MOG dark compact object has a good consistency with EHT observations. Moreover, like Table 1 and Fig. 3a, we see from Fig. 7 that increasing the STVG parameter value leads to amplify the shadow radius of the four dimensional MOG dark compact object. For $D\geq 5$ the impact of the STVG parameter in comparison with the extra dimension effect is somehow negligible and the effect of extra dimensions is dominated as it is obvious in Fig. 7
@@ VI Concluding Remarks
In this study, according to string theory, braneworld models, and AdS/CFT correspondence, we motivated to obtain a higher-dimensional MOG dark compact object solution. We investigated the behaviors of the corresponding MOG dark compact object’s shadow and deflection angle. Our main goal was to discover how extra dimensions and the STVG parameter affect the shadow of the black holes. To do this, we first developed a higher-dimensional metric for a dark compact object in STVG theory, and then provided the required general formalism to study the shadow behavior of these higher-dimensional black holes utilizing the Hamilton-Jacobi approach and Carter method to formulate the null geodesics around them and derive the corresponding effective potentials. Next, we introduced the celestial coordinates to specify the shadow shape of the higher-dimensional STVG black holes on the observer’s sky. We also derived the energy emission rate and deflection angle formulas in this higher-dimensional STVG scenario. Then, employing the constructed framework, we studied the shadow behavior, deflection angle, and energy emission rate of MOG dark compact object with extra dimensions. We computed and analyzed the significant impacts of the STVG parameter and extra dimensions on the shadow, deflection angle, and energy emission rate of the black holes within this setup. Moreover, we constrain these parameters by comparing the shadow size of M87* from EHT observations with the shadow radius of the higher-dimensional MOG dark compact object. Finally, the issue of acceleration bounds in higher-dimensional MOG dark compact object has been studied in this context in confrontation with EHT data of M87*.
@@
For the higher-dimensional MOG dark compact object, we discovered that for a fixed value of the STVG parameter $\alpha$, the shadow size of the MOG dark compact object decreases with increasing the number of extra dimensions $D$. Also, when the STVG parameter value increases with a fixed $D$, the shadow size of the MOG dark compact object increases. Also, we saw that for a fixed value of $\alpha$, the energy emission rate of the MOG dark compact object with extra dimensions extremely increases by growing $D$. Also, growing the STVG parameter increases the energy emission rate, not as mush as the extra dimensions. Therefore, we found that extra dimensions accelerate the evaporation of the MOG dark compact object with higher dimensions. Then, using the Gauss-Bonnet theorem, we have calculated the leading terms of the deflection angle in the weak-limit approximation. We have discussed the impact of the STVG parameter and the extra dimensions on this optical quantity. We noticed that for a fixed value of the STVG parameter, the deflection angle of the MOG dark compact object with extra dimensions reduces by growing the number of dimensions. Also, for a fixed value of $D$, the deflection angle of the MOG dark compact object decreases by increasing the STVG parameter $\alpha$. Furthermore, by comparing the shadow radius of the MOG dark compact object with M87* shadow released by EHT, we observed that only the shadow of four dimensional MOG dark compact object with $0.2790\leq\alpha<$ 0.279 lies in the $1$-$\sigma$ uncertainties of EHT data.
@@
In summary, we can came to conclusion that the shadow of higher-dimensional MOG dark compact object is characterized by the extra dimensions in addition to the STVG parameter. In this regard, the extra dimensions within STVG theory affect the shadows of the dark compact object by reducing their size, significantly. On the other hand, we concluded that it seems possible to detect the effects of the STVG theory via EHT. The key point here is that from Fig. 7, one can expect that the limit $\alpha<0.279$ can be apparently observed from the shadow of black holes captured by EHT. These outcomes may lead to the possibility of testing the modified theories of gravity, and their solutions and their parameters by employing astrophysical observations.

### v3

@@ III Shadow behaviour of the higher-dimensional MOG dark compact object
The gravitational field of a black hole deflects some of the light that would otherwise reach the observer, when it passes in front of it. The black hole’s apparent shape is the shadow’s boundary, which is created when photons fall into the black hole and form a dark area known as the shadow. The general formulae needed to determine the deflection angle, energy emission rate, and shadow shape for the higher-dimensional general ansatz (8)—which requires examining a test particle’s motion in spacetime—are provided in this section. Then we apply higher-dimensional MOG dark compact object line element on the formulas deduced in this section.
In MOG theory with extra dimensions, we analyze the shadow and deflection angle of the dark compact object by taking arbitrary values for the STVG parameter, $\alpha=0,0.2,0.45,0.5$, and $0.674$. In Sau:2022afl , the authors consider regular MOG dark compact objects, using nonlinear field equations to determine the effects of $\phi_{\mu}$ for $\alpha>0.674$ by adopting the ADM mass, here we examine the case for $\alpha\leq\alpha_{crit}$.
III.1 Form of the Effective potential
Initially, we aim to examine the effective potential behaviour of the dark compact object with arbitrary dimensions. To do this, the Lagrangian of the test particle:
@@ III.1 Form of the Effective potential
where $i=1,2,3,...,D-3$, and $L$ is the angular momentum and $E$ is the energy of the test particle. We apply the Carter approach to study the geodesic equations and the Hamilton-Jacobi method to analyze photon orbits around the black hole. We need to make these methods general to extra dimensions for our work. The Hamilton-Jacobi method in extra dimensions has the following form
@@
We have the following by putting the general arbitrary dimensions ansatz. (8) into Eq. (20),
@@
Since the photon is the test particle used to analyze the shadow behavior of black holes, $m$ is set to zero. Also by inserting the action (22) into Eq. (21) we derive the following equations
@@
The ”--” and ”$+$” represent respectively the ingoing and outgoing radial directions of photon motion. Also in the spacetime , we have the photon’s motion as follows which can be derived by Eqs. (26)-(29).
@@
The radial null geodesic equation is rewritten (27) and yields the effective potential,which has the following form:
@@
The black hole’s apparent shape is defined by its photons’ unstable circular orbits. The photon sphere radius $r_{0}$ corresponds to the greatest value of the effective potential at a specific distance, as shown in the equations below.
@@
Therefore, The photon sphere radius $r_{0}$ corresponds to the black hole’s maximal effective potential in spacetime (8). With arbitrary dimensions,
@@
For various values of $D$ and $\alpha$, the behavior of the effective potential for the higher dimensional MOG dark compact object is shown in Fig. 2 as a function of radial coordinate $r$. The photon sphere radius $r_{0}$, which corresponds to each value of $D$ and $\alpha$, is where the effective potential in this figure peaks. As we can observe from Fig. 2a, the higher-dimensional MOG dark compact object’s effective potential grows as $D$ increases for a constant value of $\alpha$. Furthermore, we discover from Figs. 2b and 2c that raising $\alpha$ for $D=4,6$ causes the dark compact object’s effective potential to be amplified. By examining these figures it is obvious that Higher dimensions have greater impact than $\alpha$. in the MOG dark compact object by increasing the spacetime dimensionality. Figure 2 illustrates how $D$ and $\alpha$ impact the shadow boundary of the MOG dark compact object by increasing the spacetime dimensionality. This is because the location of the black hole’s maximum effective potential, describes the black hole’s shadow boundary.
III.2 Geometric shapes of shadow
@@
The effective potential and function $\mathcal{R}$ can be rewritten in terms of impact parameters, as
@@
Putting Eq. (36) into Eq. (32) yields the following equation for two unknowns $\xi$ and $\eta$.
@@
Consequently, using Eq. (37),The photon sphere radius from Eq. (33) produces the quantity $\eta+\xi^{2}$. It is evident that the quantity $\eta+\xi^{2}$ has the dimension of the length square, while $r_{0}$ has the dimension of length.
Usually $\lambda$ and $\psi$ are used as celestial coordinates to describe the geometric shape of the shadow as perceived by the observer’s frame Vazquez:2003zm . These coordinates are
@@
We will now demonstrate the geometrical shape of the MOG dark compact object’s shadow in higher dimensions on the observer’s sky using the celestial coordinates that were presented above. To begin with, data for the black hole’s related $r_{eh},r_{0}$, and $\sqrt{\eta+\xi^{2}}$ are gathered. For the higher-dimensional MOG dark compact object, the photon sphere radius may be obtained by inserting Eq. (9) into Eq. (33). Furthermore, the radius of shadow circles for the higher-dimensional MOG dark copact object in celestial coordinates may be found by utilizing Eq. (39) and applying Eq. (9) into Eq. (37). In Table 1 we collect the numerical data associated with $r_{e}h$, and $r_{0}$ for $D=4,5,6,7,8$ and some different values of $\alpha$.
@@
Plotting the shadow circles of the MOG dark compact object with extra dimensions for various values of $\alpha$ and $D$ is possible using the information given in Table 1. Figure 3 shows the geometrical shapes of the MOG dark compact object’s shadow with extra dimensions in celestial coordinates for various values of $\alpha$ and $D$. As one can see it shows plots for constant values of $\alpha$ and illustrates how the black hole’s shadow shapes get smaller as dimension increases for a certain value of $\alpha$. Therefore, by reducing the size of shadow’s geometrical shape, the extra dimensions have a major impact on the black hole’s shadow.
@@ III.3 Energy Emission rate
Form of the energy emission rate of higher-dimensional black holes is
@@
For getting the values of the Hawking temperature of the MOG dark compact object with extra dimensions for various values of $\alpha$ and $D$, we now use the values of $r_{eh}$ from Table 1. Next, by inserting the values of $\sqrt{\eta+\xi^{2}}$ into Eq. (40), one can determine the values of $\sigma_{lim}$ for various values of $\alpha$ and $D$ which correspond to the higher-dimensional MOG dark compact object. In order to obtain the energy emission rate in terms of various values of $\alpha$ and $D$ related to the higher-dimensional MOG dark compact object, one can finally use the values of the Hawking temperature and $\sigma_{lim}$ in Eq. (41).
@@
Fig. 5 shows the energy emission rate as a function of the emission frequency for the higher-dimensional MOG dark compact object. The energy emission rate of MOG dark compact object with extra dimensions greatly increases by growing $D$, as shown in Figs. 5a and 5b, at a fixed value of $\alpha$. Consequently, we discover that in STVG with higher dimensions, extra dimensions speed up the MOG compact object’s evaporation. Also, the STVG parameter would increase the energy emission rate of MOG dark compact object, which causes it’s evaporation to accelerate.
@@ III.4 Deflection Angle
In this section a setting for examining the deflection angle of higher-dimensional black holes in the line element (8) is given. We intend to apply the Gauss-Bonnet theorem Gibbons:2008rj ; Arakida:2017hrm . To begin, we compute the optical metric on the equatorial hyperplane $\theta_{i}=\pi/2$ using the line element (8). On this hyperplane, we set $d\theta_{D-2}^{2}=d\phi^{2}$ to find
@@
Thus, following equation for calculating the deflection angle (see Refs. Gibbons:2008rj ; Arakida:2017hrm and references therein)
@@
Furthermore, the surface element of the optical metric (42) for the higher-dimensional MOG dark compact object in correspondance with the metric coefficients (9) is approximately
@@
The deflection angle of the higher-dimensional MOG dark compact object as follows
@@
The behavior of the deflection angle of the higher-dimensional MOG dark compact object is illustrated in Fig. 6 for different values of $D$ with respect to $\alpha=0.45$ in Fig. 6a and for different values of $\alpha$ with respect to $D=4,5$ in Figs. 6b and 6c, respectively. Fig. 7a, shows us that decreasing the value of the impact parameter $\xi$ results in extremely increasing the deflection angle of the MOG dark compact object. From Fig. 6a, it is obvious that for a constant value of STVG parameter, the deflection angle of the MOG dark compact object reduces by growing the number of extra dimensions. In Fig. 6b and Fig. 6c it is clear that for a fixed value of $D$,increasing STVG parameter $\alpha$ would result in decreasing the deflection angle.
IV Constraining STVG Parameter from EHT observations of M87*
@@
Figure 7 indicates the behavior of the shadow radius of the higher-dimensional MOG dark compact object in comparison to the EHT’s shadow size of M87* within $1$-$\sigma$ uncertainties as reported in Eq. (52) versus the STVG parameter $\alpha$. In this figure, it is so obvious that the shadows of the higher-dimensional MOG dark compact object associated with $D=5,6,7$ are incompatible with the observations of EHT. However, the shadow of the MOG dark compact object with $D=4$ is inside the $1$-$\sigma$ confidence level of EHT data, and in the limit $0.2790\leq\alpha<$ 0.279 the shadow radius of the four dimensional MOG dark compact object demonstrates good concordance with EHT observations. Additionally, such as in Table 1 and Fig. 3a, it is clear from Fig. 7 that increasing the STVG parameter value results in amplifying the shadow radius of the four dimensional MOG dark compact object. For $D\geq 5$ the effect of the STVG parameter in comparison to the extra dimension impact is somehow negligible and the effect of extra dimensions is dominated as it is obvious in Fig. 7
@@ VI Concluding Remarks
In this study, according to string theory, braneworld models, and AdS/CFT correspondence, we motivated to obtain a higher-dimensional MOG dark compact object solution. We investigated the behaviors of the corresponding MOG dark compact object’s shadow and deflection angle. Determining how extra dimensions and the STVG parameter affect the shadow of the black holes was our main concern. At first we developed a higher-dimensional metric for a dark compact object in STVG theory, then by using the Hamilton-Jacobi approach and Carter method to develop the null geodesics around them and calculate the corresponding effective potentials. After that, we used the celestial coordinates to determine the shadow shape of the higher-dimensional STVG black holes on the observer’s sky. In this higher-dimensional STVG framework, we also developed the formulas for the energy emission rate and deflection angle . The shadow behavior, deflection angle, and energy emission rate of a MOG dark compact object with extra dimensions were then investigated using the framework we developed. We computed and investigated the major effects of the STVG parameter and additional dimensions on the shadow, deflection angle, and energy emission rate of the black holes in this scenario. Furthermore, we restrict these parameters by comparing the shadow size of M87* obtained from EHT observations to the shadow radius of the higher-dimensional MOG dark compact object. Finally, the issue of acceleration bounds in higher-dimensional MOG dark compact object has been studied in this context in confrontation with EHT data of M87*.
@@
For the higher-dimensional MOG dark compact object, we found that for a constant value of the STVG parameter $\alpha$, the shadow size of the MOG dark compact object decreases with increasing the number of extra dimensions $D$. Also, when the STVG parameter value increases with a fixed $D$, the shadow size of the MOG dark compact object increases. Also, we saw that for a fixed value of $\alpha$, the energy emission rate of the MOG dark compact object with higher dimensions extremely goes up by increasing $D$. Furthermore, increasing the STVG parameter results in growing the energy emission rate, not as mush as the extra dimensions. Therefore, we found that extra dimensions accelerate the evaporation of the MOG dark compact object with extra dimensions. After that, utilizing the Gauss-Bonnet theorem, we determined the deflection angle’s leading terms in the weak-limit approximation. We’ve explored how the STVG parameter and extra dimensions affect this optical quantity. We noticed that for a constant value of the STVG parameter, the deflection angle of the MOG dark compact object with higher dimensions would reduce after we grow the number of dimensions. Also, for a fixed value of $D$, the deflection angle of the MOG dark compact object decreases by increasing the STVG parameter $\alpha$. Furthermore, we found that only the shadow of a four-dimensional MOG dark compact object with $0.2790\leq\alpha<$ 0.279 is inside the $1$-$\sigma$ uncertainty of EHT data.
@@
In summary, we can came to conclusion that the shadow of higher-dimensional MOG dark compact object is defined by the extra dimensions and also the STVG parameter. In this sense, the extra dimensions in STVG theory have a substantial effect on the shadows of the dark compact object, lowering their size. On the one hand, it was shown that it seems reasonable to detect the effects of the STVG theory via EHT. The crucial case here is that Fig. 7 shows that the limit $\alpha<0.279$ can be noticed from the shadows of black holes caught by EHT. These findings may be resulted in the potential of testing the modified theories of gravity, and their solutions and their parameters by utilizing astrophysical data.